OpenStudy (anonymous):

Solve the system:x + y = 5 2x – y = 1 Write your answer as an ordered pair, (x,y).

OpenStudy (anonymous):

i found x i got x=2 i need help finding y

OpenStudy (anonymous):

i think it might be y=4 is that right?

OpenStudy (anonymous):

(2,3)

OpenStudy (anonymous):

X=2 y=3

OpenStudy (anonymous):

3x = 6 x = 2 y = 3

OpenStudy (anonymous):

i got that but how did you get 3

OpenStudy (anonymous):

you plugin x

OpenStudy (anonymous):

into the equations you have

OpenStudy (anonymous):

i know that to i must have messed up somewhere can you show me your work please

OpenStudy (anonymous):

Substitute the value of x in the 1st equation which is x+y=5 So x=2, then 2+y=5 y= 5-2 y=3

OpenStudy (anonymous):

sure 2x – y = 1 4 - y = 1 4- 3 = 1 y = 3

OpenStudy (anonymous):

for the 2x-y=1 how did you get 1 when it is x+y=5?

OpenStudy (anonymous):

thank you kittykatkat14

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

it says 2x - y = 1 below the first problem

OpenStudy (anonymous):

You can substitute either problem.